For his daily training, an athlete runs around a 15 m radius track five times in 2.5 minutes. What is his average velocity in m/s?


1
Expert's answer
2021-09-08T08:13:49-0400

Given:

"R=15{\:\rm m},\quad t=2.5{\:\rm min}=150{\:\rm s},\quad n=5"


The average velocity

"v=\frac{d}{t}=\frac{2\pi R n}{t}\\=\frac{2*3.14*15*5}{150}=3.14\:\rm m/s"